PSb 4-3 Calculate Medicare Tax For each of the following employees, calculate the Medicare tax for the weekly pay period described: NOTE: For simplicity, all calculations throughout this exercise, both intermediate and final, should be rounded to two decimal places at each calculation. 1:Caleb Griffin's filing status is married filing jointly, and he has earned gross pay of $3,100. Each period he makes a 401(k) contribution of 7% of gross pay. His current year taxable earnings for Medicare tax, to date, are $248,400. Total Medicare Tax = $ 72.85 2:Anderson Fowler's filing status is single, and he has earned gross pay of $2,220 Each period he makes a 403(b) contribution of 11% of gross pay and contributes 2.5% of gross pay to a dependent care flexible spending plan. His current year taxable earnings for Medicare tax, to date, are $124,375. Total Medicare Tax = $ 52.17 3:Rick Portnoy's filing status is head of household, and he has earned gross pay of $1,300. Each period he contributes $40 to a flexible spending plan. His current year taxable earnings for Medicare tax, to date, are $206,100. Total Medicare Tax = $ 18.27 4:Shawn Buffett's filing status is married filing separately, and he has earned gross pay of $2,680. Each period he makes a 401(k) contribution of 14% of gross pay and contributes $60 to a cafeteria plan. His current year taxable earnings for Medicare tax, to date, are $198,300. Total Medicare Tax = $
Answer #1

Part 1

Medicare tax = 3100*2.35% = $72.85 (as total gross pay is above $200000 so additional 0.9% is also charged, 1.45%+0.9% = 2.35%)

Part 2

Medicare tax = 2220*1.45% = $32.19 (as total gross pay is below $200000)

Part 3

Medicare tax = (1300-40)*2.35% = $29.61 (as total gross pay is above $200000 so additional 0.9% is also charged, 1.45%+0.9% = 2.35%)

Part 4

Taxable earnings for medicare tax = 2680-60 = 2620

Total earings = 198300+2620 =$200920

Medicare tax = (2620*1.45%)+(920*0.9%) = $46.27